OpenStudy (anonymous):

Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x2 − 4x + y2 + 8y = −4

OpenStudy (anonymous):

suppose i don't want to show all my work? can i get partial credit?

OpenStudy (anonymous):

I don't want the answer. I want you to help me do it

OpenStudy (anonymous):

the idea is to complete the square twice, once for \[x^2-4x\] and again for \[y^2+8y\]

OpenStudy (anonymous):

4x^2 and 8y^2

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

have you learned how to "complete the square"?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then you can't do this problem

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i will show you how to complete the square for \(x^2-4x\) what is half of 4?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

right so we start with \[x^2-4x+y^2+8x=-4\]and turn it in to \[(x-2)^2+y^2+8y=-4+2^2\]

OpenStudy (anonymous):

or\[(x-2)^2+y^2+8y=0\] now we repeat with \(y^2+8y\)

OpenStudy (anonymous):

what is half of 8?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

so we go right to \[(x-2)^2+(y+4)^2=0+4^2\]

OpenStudy (anonymous):

or just \[(x-2)^2+(y+4)^2=16\] now your circle is in standard form and you can read off the center and radius from the equation

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I'm so sorry. when I look at numbers, they get mixed up in my head and I don't understand math in the first place.